The density of concentrated ammonia, which is 28.0% w/w NH3, is 0.899 g/mL. What volume of this reagent should be diluted to 1.0 × 103 mL to make a solution that is 0.036 M in NH3


1
Expert's answer
2022-02-20T08:40:14-0500

n1 = n2

CM = n / V

n = CM x V

n = m / M

M (NH3) = 17 g/mol

n (NH3) = 0.036 x 1 = 0.036 mol

m (NH3) = 0.036 x 17 = 0.612 g

m (NH3)28% = (0.612 x 100) / 28 = 2.186 g

Density = m / V

V = m / density

V (NH3)28% = 2.186 / 0.899 = 2.432 mL